OpenStudy (magbak):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Part 1: Find the polynomial f(x) that has the roots of –3, 5 of multiplicity 2. (4 points) Part 2: Explain how you would verify the zeros of f(x). (4 points) @Loser66 @primeralph @satellite73

OpenStudy (magbak):

@Noura11 @freddy_eighty7 @genius12 @primeralph @tcarroll010 @whpalmer4 @phi @ash2326

OpenStudy (magbak):

@satellite73

OpenStudy (magbak):

@ajprincess

OpenStudy (anonymous):

f(x) = (x+3)(x-5)^2 then expand f(-3) = 0 f(5) = 0

OpenStudy (anonymous):

taking into consideration the zero product rule, we have. \[(x - x_0)(x-x_1)(x-x_2) = 0\] where\[x_i\] are roots. given the roots we have (x +3)(x-5)(x-5). The answer is the polynomial obtained by multiplying through

OpenStudy (magbak):

Is that the answer for both parts. LOL please help me I have been studying for 14 hours straight.

OpenStudy (magbak):

@AccessDenied

OpenStudy (anonymous):

you could verify them by plugging them into the expanded polynomial. these are the values for which f(x) = 0 [the roots] plugging in any other number will not give you zero. if you don't get zero, you're polynomial or root assumptions are wrong

OpenStudy (magbak):

Ok can you sum up part one and part 2 in one post please.

OpenStudy (anonymous):

i think you'll learn a lot if you'd do that yourself :)

OpenStudy (magbak):

Come on please. I got this as my answer it is just I am so tiered and I HAVE to finish this lesson please help me out.

OpenStudy (anonymous):

\[f(x) = (x+3)(x-5)(x-5) = x^3 - 7x^2 -5x + 75\]

OpenStudy (magbak):

So that is part 1 and 2

OpenStudy (anonymous):

lol my bad. just part 1. part 2 is pretty concise above though.

OpenStudy (anonymous):

you'd plug x = -3 and x = 5 into the expanded polynomial and you should get f(x) = 0. if you don't something went wrong

OpenStudy (magbak):

Oh no problem let me sum it up.

OpenStudy (magbak):

Part 1 The polynomial is f(x)=(x+3)(x−5)(x−5)=x 3 −7x 2 −5x+75 Part 2 I plugged x = -3 and x = 5 into the expanded polynomial and I should get f(x) = 0.

OpenStudy (magbak):

@Euler271

OpenStudy (anonymous):

yes

OpenStudy (magbak):

THANK YOU THANK YOU THANK YOU THANK YOU THANK YOU THANK YOU SOOOOOOOOOOOOOOOOOOOOOO MUCH I have had a question open for 2 hours no one able to answer it THANK YOU THANK YOU I DO NOT KNOW HOW TO THANK YOU

OpenStudy (anonymous):

lol i'm glad i could help :)

OpenStudy (magbak):

Just one thing does LOL mean laugh out loud

OpenStudy (anonymous):

yes. but it isn't exactly literal lol i didn't literally laugh this time lol

OpenStudy (magbak):

Yep.

OpenStudy (magbak):

You are probably like how stupid is he.

OpenStudy (anonymous):

lol nah

OpenStudy (magbak):

Ok thank you once again. Can I tag you if I need help in the future.

OpenStudy (anonymous):

yes